OpenStudy (3psilon):

How to integrate (6-2x)^2 . (indefinite integral). I'm trying U sub but Idk what I'm doing wrong

OpenStudy (turingtest):

what did you use for u ?

OpenStudy (3psilon):

6-2x

OpenStudy (turingtest):

can you show your sork? what did you get for du? what did you get for dx?

OpenStudy (3psilon):

U = 6-2x du = -2dx \[(1/2) \int\limits_{}^{} -2U^2dx\] @TuringTest

OpenStudy (turingtest):

I'm not seeing where you get the -2 from... u=6-2x du=-2dx -> dx=-du/2

OpenStudy (3psilon):

It's to get du right ? so I multiply by 1 pretty much 2/2

OpenStudy (turingtest):

you have left your integral all mixed up with u integrated with respect to x; you can't do that. I think you are getting confused there. you need to sub in the values: u=6-2x du=-2dx -> dx=-du/2 subbing in for u and dx gives\[\int u^2(-\frac{du}2)=-\frac12\int u^2du\]

OpenStudy (3psilon):

Ohhh thank you turing!

OpenStudy (turingtest):

Welcome!

OpenStudy (3psilon):

We weren't taught how to solve for dx and plug back in but I like that way!

OpenStudy (turingtest):

I sort of never saw them do it in class that way either, I never understood why :p

OpenStudy (3psilon):

@TuringTest What answer are you getting ? I get \[\frac{-(6-2x)^{3}}{6} +c\] But my TI-89 is giving me\[\frac{4(x-3)^{3}}{3} +C\]

OpenStudy (turingtest):

it is the same answer; they appear to have factored out a -2 in your TI-89

OpenStudy (3psilon):

Ohh okay! Thanks! :)